OpenStudy (anonymous):

hich number completes the perfect square trinomial? x2 + 20x + ?

OpenStudy (anonymous):

what number squared is also 20 when added together?

OpenStudy (anonymous):

x^2+20x +C = (x+a)^2 x^2+20x+C = x^2 +2a +a^2 Compare both sides 20=2a a=10 Therefore, C=a^2=10^2=100